Main INDEX
Monthly INDEX
PREV
NEXT
User name LeRose
Log entry time 14:40:54 on April 5,2001
Entry number 59979
keyword=Shift Summary
run 3364 ended at 8:20 (last of kin 23)
Changed to kin 24
Left: p0=1.893 at 17.456 deg
Right: p0=-1.593 at 33.982 deg
Had trouble getting right bogie control screen to come up during the change. Fix was pushing right spectrometer bogie control vme reset button in the middle room.
Other crate resets were tried first and one spin off was that the right vdc high voltage went to zero (HV was reset right away but didn't notice until 10:30 that the thresholds went to zero as well)
ready at 9:00
Jean Claude did OTR tests for 20 minutes
9:22 run 3365 12C target ended 9:28
9:44 run 3366 15 cm dummy ended 9:59
10:12 run 3367 9-foil optics ended 10:26
Realized prescales hadn't been reset for runs 3365-3367
must redo (yes, I'm a dummy!)
10:26 run 3368 9-foil optics (ps=1/500/1/500/1) ended 10:40
noticed that right vdc thresholds were zero (probably the result of assorted crate resets during the kinematics change)
10:41 run 3369 still has very high deadtime (63%) with 1M events in 4 minutes
changed prescales to 1/500/12/500/1
10:47 run 3370 9-foil optics looks better (dt=7%) ended 11:01
11:08 run 3371 15-cm dummy ended 11:24
11:28 run 3372 12C ended 11:34
Change to LH2 for production running
prescales reset to production values (500/5000/35500/9500/1)
11:43 MCC says "No beam for about 10 minutes"
Problem is dump ion chamber trips.
In the end it looks like the cryotarget shifted. (See Target Down log entry)
13:45 Decision is made to warm up target for repairs.